OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (6, 2) and (7, 4)

OpenStudy (anonymous):

Should I write the same formula here or you understand it??

OpenStudy (anonymous):

I want to ask just two simple question : What is 4 - 2 = ??

OpenStudy (anonymous):

2

OpenStudy (anonymous):

And 7 - 6 = ?/

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Then just divide the two answers... \[\frac{2}{1} = ??\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

It is nothing but your slope dear...

OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = -6.9x + 5.1

OpenStudy (anonymous):

See, if you have two points that are \((x_1, y_1), (x_2, y_2)\), then: \[Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

the answer is 2 right?

OpenStudy (anonymous):

Right...